if the area of a rectangle is 192 square feet, and the ratio of width to length is 3:4, what is the width and length of the rectangle?
Answer by JulietG(1812)

Area = length * width
The question gives you the following information:
L * W = 192
W = 3/4 L
Substitute the value for W:
L * 3/4L = 192
Multiply each side by 4/3 (to get rid of the fraction)
L * L = 192 * 4/3 (256)
L squared = 256
Take the square root of each side:
L = 16
If L = 16 and W is 3/4 of that, then W = 12
.
Let's prove it by putting the values back into the original question.
L * W = 192
16 * 12 does indeed equal 192, so the values are correct.